OpenStudy (anonymous):

2x-1=x^2 -x^2+2x-1 --> (solve it with Quadratic quation to find x 7x-6=x^2 -x^2+7x-6 -> (solve with quadratic equation

OpenStudy (anonymous):

this is the quadratic equation: \[x = -b \pm \frac{\sqrt{b ^{2}-4ac}}{2a}\]

OpenStudy (anonymous):

\[-b \pm \sqrt{b ^{2}-4ac}/2a\] this is the equation for 1. a=-1; b=2; c=1.. plug in & solve

OpenStudy (anonymous):

same goes for #2

OpenStudy (anonymous):

THANK YOU

OpenStudy (anonymous):

no problem!!

OpenStudy (anonymous):

@math456 where did u get c=1 from? shouldn't it be c=-1?

OpenStudy (anonymous):

oh yeah sorry thanks for noting!!

OpenStudy (anonymous):

cool :)